Explain the effect of discretionary increase in government spending of $60 billion on the economy when the economy's marginal propensity to consume is 0.8.
Answer #1

If the marginal propensity to consume is 0.8 then the multiplier will be K = 1/ 1-MPC

= 1 / 1 - 0.8 = 1 / 0.2 = 5.

The multiplier in the economy is only 5.

An increase of the government spending of $60 billion will lead to an increase in the total GDP by 60 x 5 = $300 billion.